# Model-based Deep Learning for Joint RIS Phase Shift Compression and WMMSE Beamforming

### v2

@@ Abstract
A model-based deep learning (DL) architecture is proposed for reconfigurable intelligent surface (RIS)-assisted multi-user communications to reduce the overhead of transmitting phase shift information from the access point (AP) to the RIS controller. The phase shifts are computed at the AP, which has access to the channel state information, and then encoded into a compressed binary control message that is sent to the RIS controller for element configuration. To help reduce beamformer mismatches due to phase shift compression errors, the beamformer is updated using weighted minimum mean square error (WMMSE) based on the effective channel resulting from the actual (decompressed) RIS reflection coefficients. By unrolling the iterative WMMSE algorithm as part of the wireless communication informed DL architecture, joint phase shift compression and WMMSE beamforming can be trained end-to-end. Simulations show that accounting for phase shift compression errors during beamforming significantly improves the sum-rate performance, even when the number of control bits is lower than the number of RIS elements.
@@ I Introduction
The reconfigurable intelligent surface (RIS), a two-dimensional surface composed of several passive reflective elements, has attracted significant interest as a means to control the wireless propagation environment [1]. By jointly optimizing the RIS phase shifts and the beamformer, we can extend the signal coverage and enable cooperative signal focusing and interference cancellation along the wireless links between the access point (AP) and user equipments (UEs). This optimization takes place typically at the AP, which, having access to the channel state information (CSI), computes and transmits the phase shifts to the RIS controller.
@@
Given its importance, joint optimization of the beamformers and RIS phase shifts is an active research topic. Various methods have been proposed to tackle this problem [2, 3, 4, 5, 6]. In particular, [2] proposed an alternating optimization (AO) method to minimize the transmit power, [3] proposed a block coordinate descent (BCD) method to maximize the weighted sum-rate, [4] proposed a weighted minimum mean square error power iteration (WMMSE-PI) method with lower complexity than the previous methods, [5] proposed a deep quantization neural network (DQNN) for the optimization of discrete phase shifts for each RIS element, and [6] proposed a deep learning (DL) architecture ACFNet that compresses and decompresses the bits of the RIS phase shifts between the AP and RIS controller. With the exception of [6], most studies overlook the phase shift control process, specifically, the overhead involved in transferring information from the AP to the RIS controller.
@@
RISs are designed to be cost-effective and lack computing and sensing capabilities. As a result, computationally intensive tasks such as channel estimation [7, 8, 9, 10], and joint optimization of the beamforming and phase shifts are performed at the AP. The RIS controller, which updates the phase shifts, therefore relies entirely on the AP for phase shift reconfiguration. However, transmitting the phase shift value for each RIS element every time reconfiguration is required introduces a communication overhead that scales linearly with the number of elements, becoming a significant bottleneck in systems involving RISs with numerous elements [11]. This necessitates compression of the phase shifts which inevitably introduces decompression errors and a mismatch between the phase shifts assumed by the beamformer and those actually applied by the RIS which degrades the signal quality. Therefore, joint beamforming and phase shift optimization must account for the information transfer constraints between the AP and RIS controller.
@@
Other than the ACFNet [6], previous works do not consider the phase shift control message transmission constraints. Despite the ACFNet compressing the RIS phase shifts, this DL model does not consider the effective channel with the decompressed phase shifts when jointly optimizing the beamforming and phase shifts. Since the ACFNet does not account for this mismatch, its compression gains come at the cost of significantly decreasing the achievable sum-rate. This performance gap between conventional iterative algorithms [2, 3, 4] that were designed by analyzing the wireless communication system model, versus previous generic DL methods that try to automatically discover the model information but result in overfitting [5, 6], indicates that there is room for improvement when designing a DL architecture with the phase shift control message transmission constraints.
@@
  • We propose a joint beamforming and RIS phase shift DL architecture that compresses the phase shift information into a single control message. Compared to previous works [2, 3, 4, 5, 6], instead of individually transmitting phase shifts for each RIS element from the AP to the RIS controller, the proposed architecture takes advantage of the correlations between the phase shifts to extract and combine relevant features into the control message to reduce communication overhead.
  • Compared to conventional iterative algorithms [2, 3, 4], we jointly optimize the phase shifts and beamformer, maximizing the achievable sum-rate over multiple UEs, using end-to-end DL training. This is made possible by designing the AQE-WMMSE to use a trainable version of the iterative weighted minimum mean square error (WMMSE) algorithm based on unrolling principles [12].
  • The AQE-WMMSE accounts for the phase shift mismatch problem by explicitly integrating the compression process into the joint beamforming and phase shift optimization, and due to its model-based structure has a small hypothesis space which helps to avoid overfitting. In contrast, the generic neural networks (NNs) used in [5, 6] are prone to overfitting due to the generalized hypothesis space.
This paper is structured as follows. Section II describes the system model and the problem formulation, in Section III we propose a DL architecture and evaluate the proposed method against previous works in Section IV, finally we conclude the paper in Section V.
@@ II-A System Model
We consider the downlink of a narrowband RIS-assisted communication system, as depicted in Fig. 1, comprising an AP with $M$ transmit antennas serving $K$ UEs each with one receive antenna, and an RIS with $N$ passive phase shift elements as depicted in Fig. 1.
@@
The channels from the AP to the RIS, from the RIS to the $k$-th UE, and from the AP to the $k$-th UE are denoted by $\textbf{H}_{AR}\in\mathbb{C}^{N\times M}$, $\textbf{h}_{RU,k}\in\mathbb{C}^{1\times N}$, and $\textbf{h}_{AU,k}\in\mathbb{C}^{1\times M}$, respectively. All channels are assumed to be quasi-static and perfectly known [^1] Various techniques for channel estimation are explored in [7], [8], [9], and [10]. Assuming perfect CSI allows us to investigate upper bounds on the performance of RIS-assisted communications. at the AP. We also define the aggregate channel matrices $\textbf{H}_{RU}=[\textbf{h}_{RU,1}^{T},\ldots,\textbf{h}_{RU,K}^{T}]^{T}$, and $\textbf{H}_{AU}=[\textbf{h}_{AU,1}^{T},\ldots,\textbf{h}_{AU,K}^{T}]^{T}$. The AP beamforming vector for the $k$-th UE is denoted by $\textbf{w}_{k}\in\mathbb{C}^{M\times 1}$. Collectively, all $\textbf{w}_{k}$ satisfy a transmit power constraint $\sum_{k=1}^{K}\|\textbf{w}_{k}\|^{2}\leq P$, or $\|\textbf{W}\|^{2}_{F}\leq P$, where $\textbf{W}=[\textbf{w}_{1},\ldots,\textbf{w}_{K}]\in\mathbb{C}^{M\times K}$ is the beamforming matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm. The RIS reflection coefficients are represented by the vector $\boldsymbol{\phi}=[e^{j\theta_{1}},\ldots,e^{j\theta_{N}}]\in\mathbb{C}^{1\times N}$, where each $\theta_{n}\in[-\pi,\pi)$ denotes the phase shift applied by the $n$-th element, $1\leq n\leq N$. The phase shifts are collected in the vector $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]$.
@@
$\mathcal{R}_{k}=\text{log}_{2}\left(1+\frac{|(\textbf{h}_{AU,k}+\boldsymbol{\phi}\textbf{H}_{k})\textbf{w}_{k}|^{2}}{\sum_{l=1,l\neq k}^{K}|(\textbf{h}_{AU,l}+\boldsymbol{\phi}\textbf{H}_{l})\textbf{w}_{l}|^{2}+\sigma_{k}^{2}}\right)$ (2)
@@ II-B Problem Statement
We note that, in the ideal case when the control link is not bandwidth-limited (i.e., $B\to\infty$), problem (P1) reduces to the traditional joint beamforming and RIS phase shift optimization problem. In this scenario, $\textbf{W}_{\text{opt}}=f_{\text{w}}({\mathcal{I}})$ and $\boldsymbol{\theta}_{\text{opt}}=f_{\text{d}}(f_{\text{c}}({\mathcal{I}}))$, where $(\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}})$ denotes the optimal solution that can be obtained by AO [2], BCD [3], or WMMSE-PI [4].
@@ III Proposed DL Architecture
In previous DL works [5, 6], generic NNs are used to solve (P1). In particular, a single NN combines the functions for beamforming $f_{\text{w}}(\cdot)$ and control message $f_{\text{c}}(\cdot)$ as $(\textbf{W},\boldsymbol{\psi})=f_{\text{w,c}}(\mathcal{I})$. Then for [6], a second NN $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})$ decodes the control message $\boldsymbol{\psi}$. However, training these NNs $f_{\text{w,c}}(\cdot)$ and $f_{\text{d}}(\cdot)$ to obtain the solution $(\textbf{W},\boldsymbol{\theta})$ to (P1) can be challenging. This is because $f_{\text{w,c}}(\cdot)$, which solves for W, does not directly depend on the decoded phase shifts in $\boldsymbol{\theta}$, making the NN’s parameters have a large hypothesis space, resulting in overfitting [13, Section 5.2]. Instead, the beamforming matrix is updated indirectly with information about the phase shifts via backpropagation of a loss function $\mathcal{L}(\textbf{W},\boldsymbol{\theta})$ (see [5, eq. (5) and (13)] and [6, eq. (7) and (10)]). That is, the only way the beamforming matrix can be updated based on the phase shifts, is through the aggregate error of the batch of training samples and backpropagation of the corresponding gradients.
@@
In this paper, we propose to design the beamforming solution to explicitly depend on the phase shift values selected by the RIS controller. To avoid overfitting, we propose a wireless communication informed DL architecture AQE-WMMSE  [^2] Source code: https://github.com/AlexanderFernandes96/AQE-WMMSE to solve (P1) with two modules, as shown in Fig. 2. The first module: Auto-Quantization-Encoder (AQE), which is responsible for encoding the available information into a binary-valued control message, then decoding the control message into the RIS phase shifts. The second module: WMMSE beamforming updater, is responsible for updating the beamforming matrix with the actual RIS reflection coefficients by using the phase shifts at the output of the AQE. [^3] We note that an identical copy of the RIS controller’s AQE decoder is used at the AP to update W. Instead of using a generic NN as in [5, 6], to update the beamforming matrix, we propose a model-based approach by unrolling the WMMSE algorithm as part of our DL architecture. The input of the AQE-WMMSE is all the information available at the AP, $\mathcal{I}=\{\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}},\textbf{H}_{AU},\textbf{H}_{AR},\textbf{H}_{RU}\}$, where ($\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}})$ is the solution to (P1), obtained with the WMMSE-PI algorithm [4]. Both the WMMSE beamforming updater and the AQE modules are trained end-to-end using $\mathcal{I}$, which can be expressed in terms of solving (P1) as: $\boldsymbol{\theta}=f_{\text{d}}(f_{\text{c}}(\mathcal{I}))$ and $\textbf{W}=f_{\text{w}}(\mathcal{I};\boldsymbol{\theta})$. The notation $\textbf{W}=f_{\text{w}}(\cdot;\boldsymbol{\theta})$ signifies that W depends on $\boldsymbol{\theta}$.
@@ III-A Auto-Quantization-Encoder
Similar to an autoencoder NN, the AQE network in Fig. 2 consists of an encoder part $f_{\text{EN}}(\cdot)$ and a decoder part $f_{\text{DE}}(\cdot)$, which are both implemented by NNs. In addition, the AQE uses an intermediate quantization layer $f_{\text{Q}}(\cdot)$ between the encoder and decoder NNs [14, 15]. In relation to (P1), the NNs perform the operations $\boldsymbol{\psi}=f_{\text{c}}(\mathcal{I})=f_{\text{Q}}(f_{\text{EN}}(\mathcal{I}))$ and $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})=f_{\text{DE}}(\boldsymbol{\psi})$.
@@ III-A2 Quantizer
[…]
$\psi_{n}=Q(\tilde{\psi}_{n})=\sum_{i=1}^{D-1}a_{i}q(c_{i}(\tilde{\psi}_{n}-b_{i}))$ (4)
[…]

### v3

@@ Abstract
A model-based deep learning (DL) architecture is proposed for reconfigurable intelligent surface (RIS)-assisted multi-user communications to reduce the number of bits required for transmitting phase shift information from the access point (AP) to the RIS controller. The AP, which has access to channel state information, computes the phase shifts and encodes them into a compressed binary control message that is sent to the RIS controller for element configuration. To help reduce beamformer mismatches caused by phase shift compression errors, the beamformer is updated using the weighted minimum mean square error (WMMSE) method based on the effective channel resulting from the actual (decompressed) RIS phase shifts. By unrolling the iterative WMMSE algorithm within the wireless communication-informed DL architecture, joint phase shift compression and WMMSE beamforming can be trained end-to-end. Simulation results demonstrate that incorporating compression-aware beamforming significantly improves sum-rate performance, even when the number of control bits is lower than the number of RIS elements.
@@ I Introduction
The reconfigurable intelligent surface (RIS), a two-dimensional surface composed of several passive reflective elements, has attracted significant interest as a means to control the wireless propagation environment [Wu2021]. By jointly optimizing the RIS phase shifts and the beamformer, we can extend the signal coverage and enable cooperative signal focusing and interference cancellation along the wireless links between the access point (AP) and user equipments (UEs). This optimization typically takes place at the AP, which, having access to the channel state information (CSI), computes the phase shifts and transmits them to the RIS controller.
@@
Given its importance, joint optimization of the beamformers and RIS phase shifts is an active research topic. Various methods have been proposed to tackle this problem [Wu2019, Guo2020, Jin2024, Xu2022, Li2024]. In particular, [Wu2019] proposed an alternating optimization (AO) method to minimize the transmit power, [Guo2020] proposed a block coordinate descent (BCD) method to maximize the weighted sum-rate, [Jin2024] proposed a weighted minimum mean square error (WMMSE)-power iteration (PI) method with lower complexity than the previous methods, [Xu2022] proposed a deep quantization neural network (DQNN) for the optimization of discrete phase shifts for each RIS element, and [Li2024] proposed ACFNet, a deep learning (DL) architecture that compresses and decompresses the bits of the RIS phase shifts between the AP and RIS controller. With the exception of [Li2024], most studies overlook the phase shift control process, specifically, the binary messages involved in transferring information from the AP to the RIS controller.
@@
RISs are designed to be cost-effective and lack computing and sensing capabilities. As a result, computationally intensive tasks such as channel estimation [Yang2023, Fernandes2023, Fernandes2024, Fernandes2024a], and joint optimization of the beamforming and phase shifts are performed at the AP. The RIS controller, which updates the phase shifts, therefore relies entirely on the AP for phase shift reconfiguration. However, transmitting the phase shift value for each RIS element every time reconfiguration is required introduces a communication overhead that scales linearly with the number of elements, creating a significant bottleneck in systems involving RISs with numerous elements [Zhang2021a]. This necessitates compressing the phase shifts, which inevitably introduces decompression errors and causes a mismatch between the phase shifts assumed by the beamformer and those actually applied by the RIS, ultimately degrading the signal quality. Therefore, joint beamforming and phase shift optimization must account for the information transfer constraints between the AP and RIS controller.
@@
Except for ACFNet [Li2024], previous works do not consider the phase shift control message transmission constraints. Although ACFNet compresses the RIS phase shifts, it does not account for the effective channel with the decompressed phase shifts when jointly optimizing beamforming and phase shifts. Since ACFNet does not account for this mismatch, its compression gains come at the cost of decreasing the achievable sum-rate. This performance gap between conventional iterative algorithms [Wu2019, Guo2020, Jin2024] that were designed by analyzing the wireless communication system model, versus previous generic DL methods that try to automatically discover the model information but result in overfitting [Xu2022, Li2024], indicates that there is room for improvement when designing a DL architecture that considers phase shift control message transmission constraints.
@@
  • We propose a joint beamforming and RIS phase shift DL architecture that compresses the phase shift information into a single binary control message. Compared to previous works [Wu2019, Guo2020, Jin2024, Xu2022, Li2024], instead of individually transmitting phase shifts for each RIS element from the AP to the RIS controller, the proposed architecture takes advantage of the correlations between the phase shifts by extracting and combining relevant features into the control message to reduce the number of bits required to update the RIS.
  • Compared to conventional iterative algorithms [Wu2019, Guo2020, Jin2024], we jointly optimize the phase shifts and beamformer, maximizing the achievable sum-rate over multiple UEs, using end-to-end DL training. This is made possible by designing the AQE-WMMSE to use a trainable unrolled version of the iterative WMMSE algorithm [Monga2021a].
  • The AQE-WMMSE accounts for the phase shift mismatch by explicitly integrating the compression process into the joint beamforming and phase shift optimization. Due to its model-based structure, the hypothesis space of the model’s parameters is smaller and reduces overfitting. In contrast, the generic neural networks (NNs) used in [Xu2022, Li2024] are prone to overfitting due to the generalized hypothesis space.
This letter is structured as follows. Section II describes the system model and the problem formulation, in Section III we propose a DL architecture and evaluate the proposed method against previous works in Section IV, finally we conclude the paper in Section V.
@@ II-A System Model
We consider the downlink of a narrowband RIS-assisted communication system, as depicted in Fig. 1, comprising an AP with $M$ transmit antennas serving $K$ UEs each with one receive antenna, and an RIS with $N$ passive phase shift elements.
@@
The channels from the AP to the RIS, from the RIS to the $k$-th UE, and from the AP to the $k$-th UE are denoted by $\textbf{H}_{AR}\in\mathbb{C}^{N\times M}$, $\textbf{h}_{RU,k}\in\mathbb{C}^{1\times N}$, and $\textbf{h}_{AU,k}\in\mathbb{C}^{1\times M}$, respectively. All channels are assumed to be quasi-static and perfectly known [^1] Various techniques for channel estimation are explored in [Yang2023], [Fernandes2023], [Fernandes2024], and [Fernandes2024a]. Assuming perfect CSI allows us to investigate upper bounds on the performance of RIS-assisted communications. at the AP. We also define the aggregate channel matrices $\textbf{H}_{RU}=[\textbf{h}_{RU,1}^{T},\ldots,\textbf{h}_{RU,K}^{T}]^{T}$, and $\textbf{H}_{AU}=[\textbf{h}_{AU,1}^{T},\ldots,\textbf{h}_{AU,K}^{T}]^{T}$. The AP beamforming vector for the $k$-th UE is denoted by $\textbf{w}_{k}\in\mathbb{C}^{M\times 1}$. Collectively, all $\textbf{w}_{k}$ satisfy a transmit power constraint $\sum_{k=1}^{K}\|\textbf{w}_{k}\|^{2}\leq P$, or $\|\textbf{W}\|^{2}_{F}\leq P$, where $\textbf{W}=[\textbf{w}_{1},\ldots,\textbf{w}_{K}]\in\mathbb{C}^{M\times K}$ is the beamforming matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm. The RIS reflection coefficients are represented by the vector $\boldsymbol{\phi}=[e^{j\theta_{1}},\ldots,e^{j\theta_{N}}]\in\mathbb{C}^{1\times N}$, where each $\theta_{n}\in[-\pi,\pi)$ denotes the phase shift applied by the $n$-th element, $1\leq n\leq N$. The phase shifts are collected in the vector $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]$.
@@
$\mathcal{R}_{k}=\text{log}_{2}\left(1+\frac{|(\textbf{h}_{AU,k}+\boldsymbol{\phi}\textbf{H}_{k})\textbf{w}_{k}|^{2}}{\sum_{l=1,l\neq k}^{K}|(\textbf{h}_{AU,k}+\boldsymbol{\phi}\textbf{H}_{k})\textbf{w}_{l}|^{2}+\sigma_{k}^{2}}\right)$ (2)
@@ II-B Problem Statement
We note that, in the ideal case when the control link is not bandwidth-limited (i.e., $B\to\infty$), problem (P1) reduces to the traditional joint beamforming and RIS phase shift optimization problem. In this scenario, $\textbf{W}_{\text{opt}}=f_{\text{w}}({\mathcal{I}})$ and $\boldsymbol{\theta}_{\text{opt}}=f_{\text{d}}(f_{\text{c}}({\mathcal{I}}))$, where $(\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}})$ denotes the optimal solution that can be obtained by AO [Wu2019], BCD [Guo2020], or WMMSE-PI [Jin2024].
@@ III Proposed DL Architecture
In previous DL works [Xu2022, Li2024], generic NNs are used to solve (P1). In particular, a single NN combines the functions for beamforming $f_{\text{w}}(\cdot)$ and control message $f_{\text{c}}(\cdot)$ as $(\textbf{W},\boldsymbol{\psi})=f_{\text{w,c}}(\mathcal{I})$. Then in [Li2024], a second NN $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})$ decodes the control message $\boldsymbol{\psi}$. However, training these NNs $f_{\text{w,c}}(\cdot)$ and $f_{\text{d}}(\cdot)$ can be challenging. This is because $f_{\text{w,c}}(\cdot)$, which solves for W, is a feed-forward NN prone to overfitting due to its large hypothesis space [Goodfellow2016a, Section 5.2]. Further, the beamformer matrix does not directly depend on the decoded phase shifts in $\boldsymbol{\theta}$. Instead, W and $\boldsymbol{\theta}$ are indirectly coupled via the aggregate loss function $\mathcal{L}(\textbf{W},\boldsymbol{\theta})$ over the batch of training samples and backpropagation of the corresponding gradients (see [Xu2022, eq. (5) and (13)] and [Li2024, eq. (7) and (10)]).
@@
In this paper, we design the beamforming solution to explicitly depend on the decompressed phase shifts used by the RIS controller. To avoid overfitting, we propose a wireless communication informed DL architecture AQE-WMMSE  [^2] Source code: https://github.com/AlexanderFernandes96/AQE-WMMSE to solve (P1) with two modules, as shown in Fig. 2. The first module: Auto-Quantization-Encoder (AQE), is responsible for encoding the available information into a binary-valued control message, and then decoding the control message into the RIS phase shifts. The second module: WMMSE beamforming updater, is responsible for updating the beamforming matrix with the actual RIS reflection coefficients by using the phase shifts at the output of the AQE. [^3] We note that an identical copy of the RIS controller’s AQE decoder is used at the AP to update W. Instead of using a generic NN as in [Xu2022, Li2024], to update the beamforming matrix, we propose a model-based approach by unrolling the WMMSE algorithm as part of our DL architecture. The input of the AQE-WMMSE is all the information available at the AP, $\mathcal{I}=\{\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}},\textbf{H}_{AU},\textbf{H}_{AR},\textbf{H}_{RU}\}$, where ($\textbf{W}_{\text{opt}},\boldsymbol{\theta}_{\text{opt}})$ is the solution to (P1), obtained with the WMMSE-PI algorithm [Jin2024]. Both the WMMSE beamforming updater and the AQE modules are trained end-to-end using $\mathcal{I}$, which can be expressed in terms of solving (P1) as: $\boldsymbol{\theta}=f_{\text{d}}(f_{\text{c}}(\mathcal{I}))$ and $\textbf{W}=f_{\text{w}}(\mathcal{I};\boldsymbol{\theta})$. The notation $\textbf{W}=f_{\text{w}}(\cdot;\boldsymbol{\theta})$ signifies that the beamformer W depends on the actual RIS phase shifts $\boldsymbol{\theta}$ at the AQE output.
@@ III-A Auto-Quantization-Encoder
Similar to an autoencoder NN, the AQE network in Fig. 2 consists of an encoder part $f_{\text{EN}}(\cdot)$ and a decoder part $f_{\text{DE}}(\cdot)$, which are both implemented by NNs. In addition, the AQE uses an intermediate quantization layer $f_{\text{Q}}(\cdot)$ between the encoder and decoder NNs [Shlezinger2019, Shohat2019]. In relation to (P1), these NNs perform the operations $\boldsymbol{\psi}=f_{\text{c}}(\mathcal{I})=f_{\text{Q}}(f_{\text{EN}}(\mathcal{I}))$ and $\boldsymbol{\theta}=f_{\text{d}}(\boldsymbol{\psi})=f_{\text{DE}}(\boldsymbol{\psi})$.
@@ III-A2 Quantizer
The quantizer $\boldsymbol{\psi}=f_{\text{Q}}(\tilde{\boldsymbol{\psi}})\in\mathcal{F}=\{\textit{0},\textit{1}\}^{B}$ transforms the phase features $\tilde{\boldsymbol{\psi}}$ into the control message $\boldsymbol{\psi}$. The function $f_{\text{Q}}(\cdot)$ employs a trainable scalar quantizer $Q(\cdot)$ with $D=2^{d}$ quantization levels [Shlezinger2019]. This quantizer is applied to each of the $N_{c}$ phase features, producing a $d$-bit quantized phase feature $\psi_{n}=Q(\tilde{\psi}_{n})$ for $1\leq n\leq N_{c}$. The $N_{c}$ quantized features are then concatenated into a $B=N_{c}d$-bit control message $\boldsymbol{\psi}=[\psi_{1},\ldots,\psi_{N_{c}}]\in\mathcal{F}$.
To learn the quantization regions of $Q(\cdot)$, we adopt the quantizer structure from [Shohat2019]:
$Q(x)=\sum_{i=1}^{D-1}a_{i}q(c_{i}(x-b_{i})).$ (4)
During training, $q(\cdot)=\tanh(\cdot)$ which is a continuous function allowing for the backpropagation of gradients, while during evaluation $q(\cdot)=\text{sign}(\cdot)$ is a discontinuous step function. The trainable parameters $a_{i}$, $b_{i}$, and $c_{i}$ represent the amplitude, shift, and slope of the $\tanh$ function, respectively.